Consider the lines y = 3x + 7 and x + 2y = -10. Are the lines parallel? Explain why by comparing the slopes of the lines.
Answer by nerdybill(7384)

Convert both equations into the "slope-intercept" form:
y = mx + b
.
y = 3x + 7
is already in that form.
By inspection, the slope is 3.
.
and
x + 2y = -10
2y = -x-10
y = (-1/2)x - 5
The slope is -1/2
.
Conclusion: since the slopes are NOT the same, the lines are NOT parallel